# Beyond Memorization: Gradient Projection Enables Selective Learning in Diffusion Models

### v1

@@ Abstract
Memorization in large-scale text-to-image diffusion models poses significant security and intellectual property risks, enabling adversarial attribute extraction and the unauthorized reproduction of sensitive or proprietary features. While conventional dememorization techniques, such as regularization and data filtering, limit overfitting to specific training examples, they fail to systematically prevent the internalization of prohibited concept-level features. Simply discarding all images containing a sensitive feature wastes invaluable training data, necessitating a method for selective unlearning at the concept level.
@@
To address this, we introduce a Gradient Projection Framework designed to enforce a stringent requirement of concept-level feature exclusion. Our defense operates during backpropagation by systematically identifying and excising training signals aligned with embeddings of prohibited attributes. Specifically, we project each gradient update onto the orthogonal complement of the sensitive feature’s embedding space, thereby zeroing out its influence on the model’s weights. Our method integrates seamlessly into standard diffusion model training pipelines and complements existing defenses. We analyze our method against an adversary aiming for feature extraction. In extensive experiments, we demonstrate that our framework drastically reduces memorization while rigorously preserving generation quality and semantic fidelity. By reframing memorization control as selective learning, our approach establishes a new paradigm for IP-safe and privacy-preserving generative AI.
1 Introduction
@@
Generative vision models inevitably learn [somepalli2023diffusion, carlini2023extracting] both valuable patterns and unwanted, often proprietary, attributes from their training data. When specific visual concepts, be it a person’s likeness, a trademarked logo, or a distinctive artistic style, must never be reproduced, simply removing every image that contains them is a blunt security instrument. This approach discards rich information, significantly weakening the model’s capabilities and raising the cost of dataset curation. More critically, even if exact memorization of a training image is avoided, an adaptive adversary [ren2025reverse] can still manipulate seeds, noise inputs, or carefully crafted prompts to extract sensitive content or close approximations, revealing the model’s internalized secrets.
@@
This vulnerability to adversarial feature extraction demonstrates that conventional training-time defenses [ren2024unveiling] like regularization and data filtering, which aim to limit general overfitting, often leave residual feature traces. Similarly, inference-side techniques [gandikota2023erasing] like adversarial fine-tuning and unlearning have been repeatedly shown in empirical studies to be vulnerable to concept recurrence, making it exceedingly difficult to reliably purge a learned concept from a model’s internal feature space.
@@
What is fundamentally needed is selective learning: a robust, security-focused mechanism for models to ingest all of the beneficial content in an image while systematically blocking particular concept-level signals from ever taking root. Most prior work on memorization focuses on stopping the model from overfitting to exact training examples. Our formulation goes a critical step further: we propose a defense that blocks the model from encoding any identifiable attributes of protected content, the core features of a concept, not just the specific instance in the training set.
We offer a novel perspective by reframing the problem as one of blocking concept acquisition rather than merely pruning memorized instances. This shift transforms memorization control into a provable, concept-level security constraint. By rigorously defining and enforcing this selective learning mechanism, we offer a contribution to the field of responsible and IP-safe generative modeling.
@@
Before presenting our defense, we must first establish the security landscape: why does reducing general memorization not guarantee that models will abstain from reproducing restricted concepts? We begin by framing memorization in generative AI, identifying the training dynamics that drive models to internalize unwanted examples, and briefly surveying established techniques for curbing memorization [chen2024towards, somepalli2023understanding], underscoring why they are insufficient against concept-level extraction attacks.
@@ 1.1 Main Contributions
2 Rethinking Memorization Limits for Selective Learning
We posit that reducing generic memorization [ippolito2023preventing] by itself does not satisfy the requirements of concept-level selective learning. To understand this gap, we first analyze the mechanism and limitations of conventional memorization defenses, establishing why they fail to provide robust security guarantees against targeted concept extraction.
@@ 2.1 Background: Memorization Dynamics and Attack Surface
Memorization [arpit2017closer, schwarzschild2024rethinking, somepalli2023diffusion] refers to a model’s tendency to reproduce training examples or their close variants, creating a critical attack surface where an adversary can exploit the model’s overfitting to extract sensitive, proprietary, or private training data.
@@ 2.1.1 Definition of Memorization in Generative Models
Memorization arises when a model’s capacity and training regimen exceed what’s needed for generalization, causing high-capacity networks to act as instance-specific lookup tables.
From a learning theory perspective, memorization corresponds to low generalization performance despite a low empirical risk. The empirical risk minimization objective function
@@
If $\epsilon_{\theta}$ overfits, it learns a near instance-specific mapping from $x_{t}$ to the exact noise $\epsilon$. During sampling, the multi-step Markov chain inherent to diffusion amplifies small estimation errors, sometimes collapsing onto these memorized trajectories. Hence, the lack of explicit privacy constraints encourages the network to “remember” rather than “infer”.
@@ 2.2.1 Training-Time Defenses: Failure Modes
Techniques like exact and near-duplicate removal, controlled caption granularity, and data augmentation reduce the risk of rote reproduction [chen2024towards, somepalli2023understanding]. However, adversaries [ma2024could, balle2022reconstructing, carlini2023extracting, ma2025jailbreaking, jegorova2022survey] can sidestep these defenses by crafting prompts that exploit subtle invariances or by combining noise schedules to unearth copyrighted patches.
@@
Strategies such as early stopping, per-sample loss filtering, and Differentially Private SGD (DP-SGD) each impede over-adaptation [ren2024unveiling, chen2024exploring, chen2025enhancing, bonnaire2025diffusion, dockhorn2022differentially, zhu2018hidden]. In practice, early stopping [garnier2025early, favero2025bigger] only postpones memorization. Norm‐ and loss‐threshold detectors can be gamed by smoothing gradient signatures. DP-SGD’s [nasr2021adversary] noise injection weakens when attackers aggregate numerous model outputs, and its privacy budget frequently entails a steep utility trade-off. Ambient diffusion [daras2023ambient] reduces direct memorization but its lack of targeted control means identifiable features of copyrighted entities can persist in the latent space and later be recovered. Techniques like parameter-efficient fine-tuning and low-rank adapters (LoRA) [hu2022lora, li2023loftq, han2024parameter, hayou2024lora+] and ensemble defenses [liu2024iterative] cap the memorization budget. Despite these measures, optimization routines can still navigate memorization subspaces embedded in the weight manifold.
@@ 2.2.2 Inference-Time Strategies
Inference-time guardrails (prompt filtering, pruning, heuristic perturbations, watermarking) leave model weights unchanged, allowing adversarially engineered prompts to coax out verbatim reproductions of copyrighted content [ma2024inversion, kowalczuk2025finding, ma2024could, tsai2023ring]. Pruning [ni2025controllable] fails because memorized concepts are distributed across the network. Post-training fine-tuning and unlearning methods suffer from concept recurrence: once a concept is woven into the high-dimensional feature manifold, it is extremely difficult [zhang2024generate, suriyakumar2024unstable, george2025illusion] to surgically remove without collateral damage to generalization. Adversaries [george2025illusion] can exploit residual correlations to resurrect “forgotten” content.
@@ 2.3 Mitigating Memorization Alone Cannot Guarantee Exclusion of Restricted Concepts
3 Beyond Memorization: Gradient Projection for Selective Learning
@@
In both machine unlearning and memorization mitigation, the core mechanism is the same: steer model weights away from unwanted concepts by adjusting gradients. Unlearning methods perform this adjustment in a post-training phase to erase specific knowledge, while memorization defenses operate during training to prevent overfitting to individual samples. However, existing approaches, noise injections, extra tokens, generic regularizers, are at best heuristic. They might obscure forbidden patterns, but they offer no formal guarantee that protected features won’t leak back in under adversarial prompts or clever inference hacks. Hence, there is a pressing need for a method that explicitly prevents diffusion models from learning sensitive motifs during training.
To address the challenge of enforcing IP-safe generative modeling, we propose a gradient projection-based method that constrains model updates. Gradient projection [calamai1987projected] is a classical technique used in iterative optimization to enforce constraints by projecting the gradient onto a feasible subspace, effectively removing components that violate those constraints. This approach has been successfully applied across various domains [tian2023trainable, saha2021gradient, qiao2025gradient, hoang2024learn, huang2024unified, wang2021adversarial], including fairness in machine learning, adversarial robustness by steering updates away from vulnerable directions, and privacy-focused applications such as machine unlearning.
@@
Our method operates as follows. After computing the raw gradient update, we identify a forbidden feature subspace corresponding to copyrighted or otherwise sensitive motifs. The gradient is then orthogonally projected onto the complement of this subspace, effectively removing any aligned component. This approach adapts established projected-gradient techniques [calamai1987projected], ensuring consistent reduction of memorization potential. Implementation is simple, requiring only a set of protected embeddings, determined on-the-fly based on features that must be excluded, but demanding separate forward and backward passes for both $g_{main}$ and $g_{feat}$ which nearly doubles training time, making the method a lightweight yet theoretically robust addition to any diffusion-model training pipeline.
@@ 3.1 Method: Enforcing Concept Exclusion via Orthogonal Projection
To apply gradient projection in text‐conditioned diffusion, the first step is to isolate the visual features that carry IP leakage risk. Consider an image of Kung Fu Panda’s Po. The goal is to let the model learn the general concept of “a panda doing Kung Fu” while systematically preventing the acquisition of Po’s unique, protected likeness (e.g., small, rounded ears; yellowish-brown shorts; specific eye shape).
Since our diffusion model is text-conditioned, we use the CLIP text encoder to define the direction of the forbidden concept. We pick a primary caption $p_{\mathrm{main}}$ to capture the desired concept:
$p_{\mathrm{main}}=\text{“a panda doing Kung Fu.”}$
We first compute its training gradient using the denoising loss $\mathcal{L}$:
@@
which, unconstrained, would steer the model toward reproducing all features present in the image, including Po’s appearance.
To identify and remove the memorization directions, we introduce an auxiliary prompt $p_{\mathrm{feat}}$ that isolates a protected attribute of Po. For example:
$p_{\mathrm{feat}}=\text{“a panda wearing yellowish‐brown shorts.”}$
We compute the corresponding feature gradient:
@@
This vector $g_{\mathrm{feat}}$ defines the unwanted learning direction, the IP leakage vector.
Next, we project $g_{\mathrm{main}}$ onto the orthogonal complement of $g_{\mathrm{feat}}$. Denoting by $\varepsilon$ a small constant for numerical stability, we define the orthogonalized gradient $g_{\perp}$:
$g_{\perp}=g_{\mathrm{main}}-\lambda*\frac{\bigl\langle g_{\mathrm{main}},\,g_{\mathrm{feat}}\bigr\rangle}{\|g_{\mathrm{feat}}\|^{2}+\varepsilon}\,g_{\mathrm{feat}},$
where $\lambda$ is the scaling hyperparameter that controls the magnitude of the projected gradient. Based on the theoretical construction, setting $\lambda=1$ achieves perfect orthogonal excising. However, in practice, a lower value like $\lambda=0.1$ is required to maintain stability and optimize convergence speed, acknowledging a necessary trade-off between theoretical completeness and empirical stability.
By construction, under $\lambda=1$, $g_{\perp}$ contains no component in the direction of Po’s protected attributes.
To preserve the original update magnitude and ensure consistent learning progress, we rescale $g_{\perp}$ to match the norm of $g_{\mathrm{main}}$:
@@
Finally, the model parameters are updated using the constrained gradient:
@@
where $\eta$ is the learning rate. This procedure ensures that each update reinforces the general concept while provably preventing any increase in the model’s capacity to reproduce the copyrighted features.
Figure 2: Generation results on the “Kung Fu Panda” single-image case study using the main prompt “a panda doing kung fu” and the unlearn prompt “A huge panda wearing yellow shorts with a red and yellow waistband, with small ears, bandaged ankles, distinct fur patterns.” (a) Before finetuning, the base model produces a generic kung fu panda. (b) After naïve LoRA finetuning, the model memorizes and faithfully reproduces the copyrighted visual details from the reference image. (c) With gradient-projection finetuning, the model captures only the abstract concept of a kung fu panda without emitting any of the original image’s sensitive attributes.
@@ 3.2 Analysis: Formal Security Guarantees
At its core, the gradient projection enforces that no first-order learning signal ever flows toward the forbidden feature subspace. The following theorems underpin this robust guarantee, under $\lambda=1$.
@@ 3.2.2 Invariance of the Forbidden‐Subspace Component
3.2.3 Bounded Memorization Capacity
In diffusion training, the risk of reproducing a specific feature is proportional to the projection of the model parameters onto the feature’s gradient direction. By constraining updates to be orthogonal to $S$, we guarantee a hard cap on this risk.

### v2

# Beyond Memorization: Selective Learning for Copyright‑Safe Diffusion Model Training

@@ Abstract
Memorization in large-scale text-to-image diffusion models poses significant security and intellectual property risks, enabling adversarial attribute extraction and the unauthorized reproduction of sensitive or proprietary features. While conventional dememorization techniques, such as regularization and data filtering, limit overfitting to specific training examples, they fail to systematically prevent the internalization of prohibited concept-level features. Simply discarding all images containing a sensitive feature wastes invaluable training data, necessitating a method for selective learning at the concept level.
@@
We introduce a gradient projection method designed to enforce a stringent requirement of concept-level feature exclusion. Our defense operates during backpropagation by systematically identifying and excising training signals aligned with embeddings of prohibited attributes. Specifically, we project each gradient update onto the orthogonal complement of the sensitive feature’s embedding space, thereby zeroing out its influence on the model’s weights. Our method integrates seamlessly into standard diffusion model training pipelines and complements existing defenses. We analyze our method against an adversary aiming for feature extraction. In extensive experiments, we demonstrate that our framework drastically reduces memorization while rigorously preserving generation quality and semantic fidelity. By reframing memorization control as selective learning, our approach establishes a new paradigm for IP-safe and privacy-preserving generative AI.
1.  Introduction
@@
Generative vision models inevitably learn [41, 6] both valuable patterns and unwanted, often proprietary, attributes from their training data. When specific visual concepts, be it a person’s likeness, a trademarked logo, or a distinctive artistic style, must never be reproduced, simply removing every image that contains them is a blunt security instrument. This approach discards rich training data, significantly weakening the model’s capabilities and raising the cost of dataset curation. More critically, even if exact memorization of a training image is avoided, an adaptive adversary [38] can still manipulate seeds, noise inputs, or carefully crafted prompts to extract sensitive content or close approximations, revealing the model’s internalized secrets.
@@
This vulnerability to adversarial feature extraction demonstrates that conventional training-time defenses [37] like regularization and data filtering, which aim to limit general overfitting, often leave residual feature traces. Similarly, inference-side techniques [14] like adversarial fine-tuning and unlearning, which operate as post-hoc methods, have been repeatedly shown in empirical studies to be vulnerable to concept recurrence, making it exceedingly difficult to reliably purge a learned concept from a model’s internal feature space.
@@
What is fundamentally needed is selective learning: a robust, security-focused mechanism for models to ingest all of the beneficial content in an image while systematically blocking particular concept-level signals from ever taking root. Most prior work on memorization focuses on stopping the model from overfitting to exact training examples. Our formulation goes a critical step further: we propose a defense that blocks the model from encoding any identifiable attributes of protected content, the core features of a concept, not just the specific instance in the training set. We offer a novel perspective by reframing the problem as one of blocking concept acquisition rather than merely pruning memorized instances. This shift transforms memorization control into a provable, concept-level security constraint. By rigorously defining and enforcing this selective learning mechanism, we offer a contribution to the field of responsible and IP-safe generative modeling.
@@
Before presenting our defense, we must first establish the security landscape: why does reducing general memorization not guarantee that models will abstain from reproducing restricted concepts? We begin by framing memorization in generative AI, identifying the training dynamics that drive models to internalize unwanted examples, and briefly surveying established techniques for curbing memorization [9, 42], underscoring why they are insufficient against concept-level extraction attacks.
@@ 1.1  Main Contributions
2.  Rethinking Memorization Limits for Selective Learning
We posit that reducing generic memorization [23] by itself does not satisfy the requirements of concept-level selective learning. To understand this gap, we first analyze the mechanism and limitations of conventional memorization defenses, establishing why they fail to provide robust security guarantees against targeted concept extraction.
@@ 2.1  Background: Memorization Dynamics and Attack Surface
Memorization [1, 40, 41] refers to a model’s tendency to reproduce training examples or their close variants, creating a critical attack surface where an adversary can exploit the model’s overfitting to extract sensitive, proprietary, or private training data.
@@ 2.1.1 Definition of Memorization in Generative Models
Memorization arises when a model’s capacity and training regimen exceed what’s needed for generalization, causing high-capacity networks to act as instance-specific lookup tables. From a learning theory perspective, memorization corresponds to low generalization performance despite a low empirical risk. The empirical risk minimization objective function
@@
If $\epsilon_{\theta}$ overfits, it learns a near instance-specific mapping from $x_{t}$ to the exact noise $\epsilon$. This effectively embeds training examples in the learned score field. During sampling, the iterative reverse‑diffusion process repeatedly queries this biased score estimate, causing trajectories to drift toward these memorized regions of the data manifold. In the absence of explicit privacy constraints, the model is therefore incentivized to “remember” training examples rather than learn a generalizable score function.
@@ 2.2.1 Training-Time Defenses: Failure Modes
Techniques like exact and near-duplicate removal, controlled caption granularity, and data augmentation reduce the risk of rote reproduction [9, 42]. However, adversaries [31, 2, 6, 29, 25] can sidestep these defenses by crafting prompts that exploit subtle invariances or by combining noise schedules to find copyrighted patches.
@@
Strategies such as early stopping, per-sample loss filtering, and Differentially Private SGD (DP-SGD) each impede over-adaptation [37, 7, 8, 4, 11, 55]. In practice, early stopping [15, 13] only postpones memorization. Norm‐ and loss‐threshold detectors can be gamed by smoothing gradient signatures. Although DP‑SGD [33] provides formal privacy guarantees, achieving a meaningful privacy budget often requires injecting substantial noise during training, which can significantly degrade model utility. Moreover, if the privacy parameters are set too loosely, the resulting model may still leak information that attackers can exploit, potentially by aggregating many model outputs, even though the DP guarantee itself does not weaken under such aggregation. Ambient diffusion [10] reduces direct memorization but its lack of targeted control means identifiable features of copyrighted entities can persist in the latent space and later be recovered. Techniques like parameter-efficient fine-tuning and low-rank adapters (LoRA) [21, 27, 17, 18] and ensemble defenses [28] cap the memorization budget. Despite these measures, optimization routines can still navigate memorization subspaces embedded in the weight manifold.
@@ 2.2.2 Inference-Time Strategies
Inference‑time, post‑hoc guardrails, such as prompt filtering, output pruning, heuristic perturbations, or watermarking, operate only at the interface level and leave the model weights untouched. Because the underlying generative distribution remains unchanged, adversarially crafted prompts can still navigate the model’s latent space and extract memorized or near‑verbatim copyrighted content [30, 26, 31, 47]. Pruning [34] fails because memorized concepts are distributed across the network. Post-training fine-tuning and unlearning methods suffer from concept recurrence: once a concept is woven into the high-dimensional feature manifold, it is extremely difficult [53, 43, 16] to surgically remove without collateral damage to generalization. Adversaries [16] can exploit residual correlations to resurrect “forgotten” content.
@@ 2.3  Mitigating Memorization Alone Cannot Guarantee Exclusion of Restricted Concepts
3.  Beyond Memorization: Gradient Projection for Selective Learning
@@
The key mechanism in selective learning is to modify gradient updates so that training steers the model away from unwanted concepts. Unlearning methods [50, 37] perform this adjustment in a post-hoc phase to erase specific knowledge, while memorization defenses [42, 13] operate during training to prevent overfitting to individual samples. However, these approaches, including, noise injections, extra tokens, generic regularizers, are at best heuristic. They might obscure forbidden patterns, but they offer no formal guarantee that protected features won’t leak back in under adversarial prompts or clever inference hacks. Hence, there is a pressing need for a method that explicitly prevents diffusion models from learning sensitive motifs during training.
To address the challenge of enforcing IP-safe generative modeling, we propose a gradient projection-based method. Gradient projection [5] is a classical technique used in iterative optimization to enforce constraints by projecting the gradient onto a feasible subspace, effectively removing components that violate those constraints. This approach has been successfully applied across various domains [45, 39, 36, 20, 22, 48], including fairness in machine learning, adversarial robustness by steering updates away from vulnerable directions, and privacy-focused applications such as machine unlearning.
@@
Our method operates as follows. After computing the raw gradient update, we identify a forbidden feature subspace corresponding to copyrighted or otherwise sensitive motifs. The gradient is then orthogonally projected onto the complement of this subspace, effectively removing any aligned component. This approach adapts established projected-gradient techniques [5], ensuring consistent reduction of memorization potential making our method a robust addition to any diffusion-model training pipeline.
@@ 3.1  Method: Enforcing Concept Exclusion via Orthogonal Projection
To apply gradient projection in a text-conditioned diffusion model, we first isolate the visual features that pose an IP-leakage risk. Consider an image of Kung Fu Panda’s Po. Our objective is to let the model learn the general concept of “a panda doing Kung Fu” while preventing it from acquiring Po’s distinctive, protected likeness (e.g., rounded ear shape, yellowish-brown shorts, characteristic facial features).
Because the model is text-conditioned, we use the CLIP text encoder to define directions in concept space. We begin by selecting a primary caption $p_{\mathrm{main}}$ that captures the desired concept:
$p_{\mathrm{main}}=\text{``a panda doing Kung Fu.''}$
Using the denoising loss $\mathcal{L}$, we compute the corresponding training gradient:
@@
which, if left unmodified, would encourage the model to learn all features present in the training image, including Po’s copyrighted appearance.
To identify the directions responsible for memorizing protected attributes, we introduce an auxiliary prompt $p_{\mathrm{feat}}$ that isolates a specific copyrighted feature:
$p_{\mathrm{feat}}=\text{``a panda wearing yellowish-brown shorts.''}$
The gradient associated with this prompt,
@@
serves as the IP-leakage direction we aim to remove. We then project $g_{\mathrm{main}}$ onto the orthogonal complement of $g_{\mathrm{feat}}$. With $\varepsilon$ added for numerical stability, the orthogonalized gradient is
$g_{\perp}=g_{\mathrm{main}}-\lambda\,\frac{\langle g_{\mathrm{main}},\,g_{\mathrm{feat}}\rangle}{\|g_{\mathrm{feat}}\|^{2}+\varepsilon}\,g_{\mathrm{feat}}.$
The hyperparameter $\lambda$ controls the strength of the projection. Under $\lambda=1$, the resulting $g_{\perp}$ contains no component aligned with Po’s protected attributes. To maintain consistent learning dynamics, we rescale the orthogonalized gradient to match the norm of the original update:
@@
Finally, we update the model parameters using this constrained gradient:
@@
where $\eta$ is the learning rate. This ensures that each step reinforces the intended concept while provably preventing the model from increasing its capacity to reproduce copyrighted features.
Figure 2: Generation results on the “Kung Fu Panda” single-image case study using the main prompt “a panda doing kung fu” and the forbidden prompt “A huge panda wearing yellow shorts with a red and yellow waistband, with small ears, bandaged ankles, distinct fur patterns.” (a) Before finetuning, the base model produces a generic kung fu panda. (b) After naïve LoRA finetuning, the model memorizes and faithfully reproduces the copyrighted visual details from the reference image. (c) With gradient-projection finetuning, the model captures only the abstract concept of a kung fu panda without emitting any of the original image’s sensitive attributes.
@@ 3.2  Analysis: Formal Security Guarantees
The following theorems formalize this guarantee: when $\lambda=1$, the projected gradient is exactly orthogonal to the forbidden feature subspace, removing all first‑order learning signals along those directions in each constrained update.
@@ 3.2.1 Zero First‐Order Improvement on Forbidden Features
The first-order constraint in Theorem 3.1 is applied iteratively across all $T$ updates. By enforcing $\langle g_{proj},g_{feat}\rangle=0$ at each step, we ensure the parameter state $\theta_{T}$ remains within the orthogonal complement of the forbidden subspace $S_{f}$ relative to the initialization $\theta_{0}$.
@@ 3.2.2 Invariance of the Forbidden‐Subspace Component
3.2.3 Bounded Memorization Capacity: Geometric Interpretation
In diffusion training, the risk of reproducing a specific feature can be approximated by the projection of the parameter vector onto the feature’s gradient direction. Under a first-order (NTK) approximation, we define the memorization capacity $M_{f}(\theta)$ of a forbidden feature $f$ as
$M_{f}(\theta)\;=\;\|P_{S_{f}}\theta\|^{2},\qquad S_{f}=\mathrm{span}\{g_{\mathrm{feat}}\},$
where $P_{S_{f}}$ denotes the orthogonal projector onto the forbidden feature subspace.
Geometrically, $M_{f}(\theta)$ measures the alignment of the parameter state with the feature’s Fisher-sensitive direction. In the Neural Tangent Kernel (NTK) view [24, 46], this alignment quantifies the model’s directional sensitivity to the protected attribute.
By constraining each update $\Delta\theta$ to lie in $S_{f}^{\perp}$ (i.e., $P_{S_{f}}\,\Delta\theta=0$), we ensure that the projection onto $S_{f}$ does not increase:
$P_{S_{f}}(\theta_{t+1})\;=\;P_{S_{f}}(\theta_{t}).$
Thus, the model’s first-order capacity to encode or reconstruct the forbidden feature remains bounded by its initial value. This provides a structural safeguard against feature memorization that is independent of standard overfitting metrics.